OpenStudy (anonymous):

A company that makes nails finds that it would sell 42 boxes of nails per month if they are priced at $20 per box. If the company lowers the price to $10 per box, then more people would purchase nails and the company could expect to sell 52 boxes per month. Find the equation of the line for the demand equation if x represents the number of boxes sold and y is the price of a box of nails.

OpenStudy (anonymous):

y = 62 + x y = 42 − x y = 42 + x y = 62 − x

OpenStudy (igreen):

Sorry, I'm not sure..

geerky42 (geerky42):

hmm wait silly me, i misread question

geerky42 (geerky42):

hmm, sorry, this question confused me like iGreen

OpenStudy (mathmale):

It appears that you're given two different points on the same (demand) "curve:" (42,$20) and (52, $10), where 42 and 52 are x-values (number of boxes sold), and $10 and $20 are unit prices for the nails. Try applying the slope formula to find the slope of the line through these two points. Its slope should be negative. Once you have that slope, you can determine the equation of the demand function by substituting your slope (m) and either (42,$20) or (52, $10) into the point-slope formula for a straight line.

OpenStudy (anonymous):

Did you get A??

OpenStudy (mathmale):

I'm glad to help with the problem solving, but I do not deal in A, B, C or D. Want to see your work and to understand your reasoning.

OpenStudy (igreen):

Slope Formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (mathmale):

Hint: Going through the procedure I've outlined, you want to solve the resulting equation for y, because all of the answer choices begin with y.

OpenStudy (igreen):

Right..but how is finding the slope supposed to help?

geerky42 (geerky42):

we are asked to find equation from given point (42,$20) and (52, $10) So we would need to use this formula: \(y-y_1=m(x-x_1)\) hence slope formula should be used